VIBRATIONAL FUNDAMENTALS OF GeC_{n} CHAINS TRAPPED IN SOLID ARGON

\underline{W. R. M. GRAHAM}, D. L. ROBBINS, and C. M. L. RITTBY, Department of Physics and Astronomy,Texas Christian University, Fort Worth, TX 76129.

Experimental and theoretical studies on the structures and vibrational fundamentals of a novel family of germaniumcarbon clusters (Ge_{n}C_{m}) that were initiated with our earlier identification of the GeC_{3}Ge cluster ^{a} have now been extended to include GeC_{n} chains. The new clusters, which are formed by the Nd-YAG laser ablation of germanium and graphite and are trapped in solid Ar at 10 K, have been identified using Fourier transform infrared (FTIR) measurements coupled with density functional theory calculations (DFT). The assignment of carbon stretching fundamentals for GeC_{4}, GeC_{7}, and Ge_{9} will be discussed. FTIR measurements of ^{1}^{3}C isotopic shifts for the assignments are in good agreement with the DFT predictions.

aD. L. Robbins, C. M. L. Rittby, and W. R. M. Graham, J. Chem. Phys. 114, 3570 (2001).